a) What is the wavelength of the carrier wave of a campus radio station, broadcasting at a frequency of 97.2 MHz (million cycles per second or million hertz)?


b) What is the frequency of a red laser beam, with a wavelength of 670 nm, which your astronomy instructor might use to point to slides during a lecture on galaxies?



c) Can you hear sounds in space? Why or why not?


1
Expert's answer
2021-02-28T07:22:42-0500

(a)

"\lambda=\dfrac{c}{f}=\dfrac{3\cdot10^8\ \dfrac{m}{s}}{97.2\cdot10^6\ Hz}=3.08\ m."

(b)

"f=\dfrac{c}{\lambda}=\dfrac{3\cdot10^8\ \dfrac{m}{s}}{670\cdot10^{-9}\ m}=448\cdot10^{12}\ Hz=448\ THz."

(c) The sounds can't be heard in space. Because the sound wave can't propagate in vacuum since there are very few particles to vibrate, bump each other and propagate that wave.